OpenStudy (anonymous):

x^7+x^6 -------- x I think the answer is 1^7+ 1^6.... am I right??

OpenStudy (anonymous):

\[\frac{x^7+x^6}{x}\] can you factor out an x from both pieces on top? \[\frac{(x)x^6+(x)x^5}{x}\to \frac{(x)(x^6+x^5)}{x}\]

OpenStudy (anonymous):

so I'm supposed to multiply the two x's could you walk me through it...I have an example with numbers which I found very easy but the letters have me thrown off

OpenStudy (anonymous):

No this lets you do the division easier. \[\frac{(x)(x^6+x^5)}{x}\to \left(\frac{x}{x}\right)\frac{x^6+x^5}{1}\] you know what (x/x) is?

OpenStudy (anonymous):

I thought x/x=1

OpenStudy (anonymous):

Yes. so it cancels and what are you left with?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

please forgive me if im wrong I'm completely illiterate when it comes to math

OpenStudy (anonymous):

you are left with \[x^6+x^5; x\neq 0\]

OpenStudy (anonymous):

ok I think I understand that, if I can make my brain understand numbers with letters this may become easy for me...thanks

OpenStudy (anonymous):

Keep trying:) variables are more general.

OpenStudy (anonymous):

btw can you tell me what the line through the equal sign means...I know it's not equal to but is that all of it?

OpenStudy (anonymous):

yes. its basically a short hand way of saying that x can't be zero since in the original equation you are dividing by x. And you cannot divide by 0

OpenStudy (anonymous):

okay thanx again!!!